The sum of four times a number and 4 more is the same as the difference between -31 and twice the number. What is that number?
Answer by math_helper(2461)


Let x=the number
4x+4 (sum of four times a number and 4 more)
-31-2x (difference between -31 and twice the number)
Set the above two things equal to each other ("…the same as…")
4x+4 = -31-2x
6x+4 = -31 (added 2x to both sides)
6x = -35 (subtracted 4 from both sides)
Ans: x = -35/6
Check:
4(-35/6) + 4 = (-140+24)/6 = -116/6
-31-2(-35/6) = (-186+70)/6 = -116/6 (ok)